# LLM-Enhanced User-Item Interactions: Leveraging Edge Information for Optimized Recommendations

### v1

@@ 3.3. Pre-training Graph Attentive LLM with Crowd Contextual Prompts
Pre-training is to initially train the of our graph attentive GPT-2 model on a larger corpus of recommendation text data, including data collection, tokenization, model architecture, pre-training objective, and optimization procedure.
@@
We first collect a large textual corpus from diverse sources: user descriptions, item descriptions, user reviews for items, historical events that users interact (e.g., rate or purchase) with items, in order to ensure that the graph attentive LLM model learns robust representations of languages.
@@ 3.3.2. The Structure of Crow Contextual Prompts
Figure  2 show that we define the unique structure of our crow contextual prompts for pre-training a LLM recommender.
User/Item Tokens. Our prompts include two unique tokens: user tokens and item tokens. User and item tokens are the targeted entities in a recommender system. In other words, the representation learning and generative recommendation tasks are centered around users and items. As a result, we add userID and itemID tokens into the corpus vocabulary. However, we ensure that userID and itemID tokens are not tokenized and decomposed into smaller token units. We expect the embedding of userID and itemID can embed rich semantic information about users, such as user profiles, demographics, reviews, and preferences, information propagated from items, and rich semantic information about items, such as item descriptions, item functionalities, item reviews, information propagated from users.
  • User and/or Item Contents. Aside from userIDs and itemIDs, we use the content (attributes) information of users and or items to develop content-related prompts, such as, the title of an item is ¡TITLE¿, the brand of an item is ¡BRAND¿, the product categories of an item are ¡CATEGORIES¿. the description of an item is ¡DESCRIPTIONS¿. In our experiments, user contents are removed due to privacy concerns. We only have item contents.
@@
  • User-Item Interaction (e.g., Purchase) Events . Finally, we incorporate the purchase events as prompts, such as: a user has interacted with (purchased) ¡ITEM LIST¿.
In this way, we aim to augment the prompt texts and enrich the contextual environment that simulates the preference, characteristics, functionality, categorization, opinion, and first-order and second-order social or network dimensions of a real recommender system in a language modality. These crowd contextual prompts are used as training data to pre-train our graph attentive LLM.
Figure 2. Crowd Contextual Prompts based on Recommendation Data.
@@ 3.3.3. The Optimization Objective and Procedures of Pre-training
(5) $L_{\text{pre-train}}=-\sum_{i}\log P(T_{i+1}|T_{i},T_{i-1},\ldots,T_{1};\Theta),$
where $T_{i},T_{i-1},\ldots,T_{1}$ is the first $i$ tokens, and $\Theta$ indicates the weights of the LLM, the objective is to predict the $i+1$ token $T_{i+1}$. By solving the optimization objective, the graph attentive LLM can combine knowledge from different information sources to learn more comprehensive portraits of users and items. By integrating graph structured attention, the graph attentive LLM can model the first-order and second-order edge connectivity among users and items.
@@ 3.4. Fine-tuning Graph Attentive LLM with Personalized Predictive Prompts
After the pre-training step, the graph attentive LLM learns the contextual knowledge of users, items, relationships of a recommender system’s world. However, the optimization objective in the pre-training step is focusing on maximizing textual generation accuracy in a language sequence, instead of recommending personalized items. Therefore, in the fine-tuning step, we develop 1) personalized predictive prompts and 2) recommendation loss functions of fine-tuning to incentivize the graph attentive LLM to shift model focuses from text generation to generating accurate personalized item recommendations.
@@ 3.4.1. Personalized Predictive Prompts
When fine-tuning the pre-trained graph attentive LLM, we introduce the personalized predictive Prompt method. Our idea is to use the historical purchase events of users for items as prompts to guide the LLM to learn user preferences for items. Figure 3 shows that we convert a user’s interaction (e.g., rating, purchases) history with all items into past tense texts, combined with a future tense trigger (e.g., user $i$ will purchase ????), to motivate the graph attentive LLM to generate item recommendations for a user.
Figure 3. A Personalized Predictive Prompt.
@@ 3.4.2. The Optimization Objective and Training Procedure of Fine-tuning
Different from learning the world knowledge of a recommendation system, the fine-tuning stage is to adapt the pre-trained LLM to personalized item recommendations. During the fine-tuning process, we will integrate personalized predictive prompts into LLM so that the model can model the historical item purchase event of a user to generate a list of items $R_{i}$ to the specific user, based on the prompt. The generative recommended items are then compared with the actual purchase records of the specific user. The resulting loss function derived from this comparison is utilized as the optimization objective of fine-tuning. In particular, we define the generative probability that measures whether LLM recommendations are statistically close to historical user purchase records, which is given by Equation 6:
(6) $L_{\text{fine-tune}}=-\sum_{i}\log P(R_{i}|Pr_{i},\Theta),$
where $Pr_{i}$ is the prompt for fine-tuning stage, and $R_{i}$ is the final recommendation list, $\Theta$ denotes the LLM weights. In summary, fine-tuning is to optimize the recommendation loss.
@@ 3.5. Graph Attentive LLM for Item Recommendations in Deployment
After pre-training and fine-tuning, given a testing user $i$, we convert the user’s interaction records with items into a personalized predictive prompt as the input of the graph attentive LLM. Therefore, the classic recommendation engine in production can be viewed as a wrapper, where the request is reconstructed as a prompt - the same as the prompt in the fine-tuning stage, and the graph attentive LLM will provide a recommendation score for each user-item pair.
In order to reduce the serving latency in production and alleviate the peak load pressure, the proposed graph attentive LLM model can be deployed onto both offline and online GPU clusters. The offline pipeline focuses on the batch processing and calculates the relevance between a user and the candidate items to recommend, shown as $\mathbf{LLM}(X_{ij};\Theta)$, where $X_{ij}$ is the user-item interaction prompt of user $i$ and item $j$, and $\Theta$ denotes the weights for the $\mathbf{LLM}$ backbone. The batch processing can directly be applied in offline recommendation scenarios such as promotional emails and notifications. The results can also be used for warming up online cache to minimize redundant computations.
In a typical online recommendation scenario, the prompt containing user, item and the interaction information is sent to the graph attentive LLM model, and the top $M$ items with the highest scores are selected as recommendations by comparing the probability scores against each other,
@@
In this case, the proposed method can easily be integrated into most common recommender systems in industry. The additional pressure caused by GPU serving can be handled by offline (batch) precomputation and online caching warm-up.
@@ 4. Experimental Results
We conduct empirical experiments to answer the following questions: 1) can our method generate more accurate recommendations? 2) what are the contributions of different technical components? 3) what are the contributions of second order relationship and item background information? 4) what are the impacts of different attention mechanisms? 5) parameter sensitivity and robustness.
@@ 4.1.1. Data Description
We used seven public recommendation datasets: Amazon (AM)-Beauty dataset, AM-Toys dataset, AM-Sports, AM-Luxury, AM-Scientific, AM-Instruments dataset (McAuley and Yang, 2016). We binarized the user-item interaction matrix by scores. If the score is greater than 3, there is a connection between a user and an item. For each user in the dataset, we randomly select 80% of interactions for training, 10% for validation, and 10% for testing, with at least one sample selected in both the validation and test sets. According to the prompt construction method in Section  3.3, we constructed the data for pretraining. Table  1 shows The main dataset statistics.
@@ 4.1.3. Baseline Algorithms
We compared our method with various approaches including ID-based and Attention-based methods:
  • Multi-VAE (Liang et al., 2018) is an ID-based collaborative filtering method that completes recommendation tasks by using a polynomial likelihood variational autoencoder to reconstruct ratings.
@@
  • $S^{3}Rec$ (Zhou et al., 2020) extends BERT4Rec by adding auxiliary tasks such as item attribute prediction to enhance MLM, which can integrate content features for self-supervised learning.
@@ 4.1.4. Hyper parameters and Settings
We conducted experiments using GPT-2 as the base model. We set the maximum input length to 1024, the token embedding dimension to 768, and the vocabulary length of natural language tokens to 50257. In the pre-training stage, we first trained 10 epochs using crowd contextual data to optimize LLM and then trained 100 rounds using user-item interaction data. In the fine-tuning stage, we used 50 epochs for recommendation-oriented fine-tuning of LLM.
@@ 4.1.5. Experimental Environment
All experiments were conducted on Ubuntu 22.04.3 LTS OS, Intel(R) Core(TM) i9-13900KF CPU, with the framework of Python 3.11.5 and PyTorch 2.0.1.
@@ 4.2.1. Overall Comparison
This experiment aims to answer: Can our model really generate more accurate recommendation results through the natural language processing method? We compared our model with several baseline models on various Amazon datasets. The baseline models used for comparison include ID-based and Attention-based methods. Our model was tested on the same dataset as these baseline models to ensure fairness and accuracy in the comparison. The experimental results are shown in Table  2, and our model performs well in seven Amazon datasets. Recall@20, Recall@40, and NDCG@100 are superior to the baseline models. This indicates that LLMs have strong capabilities in understanding text, and capturing user preferences and needs, thereby promoting the accuracy of recommendations. Overall, the experimental results support our hypothesis that our model can generate more accurate recommendations through the graph attentive LLM. This discovery is important for research and the practical application of recommendation systems.
@@
Dataset Metric Multi-VAE MD-CVAE BERT4Rec $S^{3}$ec UniSRec FDSA SASRec GRU4Rec LLM- NoPretrain LLM- NoFineTune LLM- NoGKIA LLM- NoGHIP Ours
AM-Beauty Recall@20 0.1295 0.1472 0.1126 0.1354 0.1462 0.1447 0.1546 0.0997 0.0464 0.0441 0.1225 0.1267 0.1590
Recall@40 0.1720 0.2058 0.1677 0.1789 0.1898 0.1875 0.2071 0.1528 0.0709 0.0691 0.1665 0.1799 0.2177
NDCG@100 0.0835 0.0835 0.0781 0.0867 0.0907 0.0834 0.0949 0.0749 0.0339 0.0323 0.0790 0.0827 0.1029
AM-Toys Recall@20 0.1076 0.1291 0.0853 0.1064 0.1110 0.0972 0.0869 0.0657 0.0477 0.0580 0.0896 0.0858 0.1349
Recall@40 0.1558 0.1804 0.1375 0.1524 0.1457 0.1268 0.1146 0.0917 0.0689 0.1003 0.1272 0.1179 0.1873
NDCG@100 0.0781 0.0844 0.0532 0.0665 0.0638 0.0662 0.0525 0.0439 0.0330 0.0481 0.0612 0.0594 0.0876
AM-Sports Recall@20 0.0659 0.0714 0.0521 0.0616 0.0714 0.0681 0.0541 0.0720 0.0449 0.0394 0.0555 0.0558 0.0764
Recall@40 0.0975 0.1180 0.0701 0.0813 0.1143 0.0866 0.0739 0.1086 0.0719 0.0613 0.0846 0.0830 0.1240
NDCG@100 0.0446 0.0514 0.0305 0.0438 0.0504 0.0475 0.0361 0.0498 0.0322 0.0278 0.0391 0.0379 0.0535
AM-Luxury Recall@20 0.2306 0.2771 0.2076 0.2241 0.3091 0.2759 0.2550 0.2126 0.1872 0.1885 0.2474 0.2679 0.3066
Recall@40 0.2724 0.3206 0.2404 0.2672 0.3675 0.3176 0.3008 0.2522 0.2233 0.2254 0.2880 0.3028 0.3441
NDCG@100 0.1697 0.2064 0.1617 0.1542 0.2010 0.2107 0.1965 0.1623 0.1223 0.1235 0.1834 0.2065 0.2331
AM-Scientific Recall@20 0.1069 0.1389 0.0871 0.1089 0.1492 0.1188 0.1298 0.0849 0.0708 0.0668 0.1383 0.1206 0.1480
Recall@40 0.1483 0.1842 0.1160 0.1541 0.1954 0.1547 0.1776 0.1204 0.1037 0.0960 0.1822 0.1575 0.1908
NDCG@100 0.0766 0.0872 0.0606 0.0715 0.1056 0.0846 0.0864 0.0594 0.0568 0.0465 0.0940 0.0810 0.1072
AM-Instruments Recall@20 0.1096 0.1398 0.1183 0.1352 0.1684 0.1382 0.1483 0.1271 0.0766 0.0727 0.1387 0.1426 0.1698
Recall@40 0.1628 0.1743 0.1531 0.1767 0.2239 0.1787 0.1935 0.1660 0.1004 0.0948 0.1741 0.1779 0.2265
NDCG@100 0.0735 0.1040 0.0922 0.0894 0.1075 0.1080 0.0934 0.0998 0.0500 0.0478 0.1042 0.1044 0.1312
AM-Food Recall@20 0.1062 0.1170 0.1036 0.1157 0.1423 0.1099 0.1171 0.1140 0.0224 0.0204 0.1275 0.1264 0.1438
Recall@40 0.1317 0.1431 0.1284 0.1456 0.1661 0.1317 0.1404 0.1389 0.0299 0.0274 0.1559 0.1487 0.1673
NDCG@100 0.0727 0.0863 0.0835 0.0926 0.1024 0.0904 0.0942 0.0910 0.0153 0.0141 0.0898 0.0963 0.1119

### v2

@@ 3.3. Pre-training Graph Attentive LLM with Crowd Contextual Prompts
Pre-training is to initially train our graph-attentive GPT-2 model on a larger corpus of recommendation text data, including data collection, tokenization, model architecture, pre-training objective, and optimization procedure.
@@
We first collect a large textual corpus from diverse sources: user descriptions, item descriptions, user reviews for items, and historical events that users interact (e.g., rate or purchase) with items, to ensure that the graph-attentive LLM model learns robust representations of languages.
@@ 3.3.2. The Structure of Crow Contextual Prompts
Fig. 2 shows that we define the unique structure of our crow contextual prompts for pre-training an LLM recommender.
User/Item Tokens. Our prompts include two unique tokens: user tokens and item tokens. User and item tokens are the targeted entities in a recommender system. In other words, the representation learning and generative recommendation tasks are centered around users and items. As a result, we add userID and itemID tokens into the corpus vocabulary to reflect user and item information. We expect the embedding of the user ID and item ID to be able to embed rich semantic information about users, such as user profiles, demographics, reviews, and preferences, information propagated from items, and rich semantic information about items, such as item descriptions, item functionalities, item reviews, information propagated from users.
To keep these special tokens, we revise the original tokenizer to prevent it from decomposing them into smaller ones. If they were broken into smaller pieces, they would have had difficulty precisely representing rich semantic information for users and items.
  • User and/or Item Contents. Aside from user IDs and item IDs, we use the content (attributes) information of users and items to develop content-related prompts, such as, the title of an item is ¡TITLE¿, the brand of an item is ¡BRAND¿, and the product categories of an item are ¡CATEGORIES¿. The description of an item is ¡DESCRIPTIONS¿. In our experiments, user contents are removed due to privacy concerns. We only have item contents.
@@
  • User-Item Interaction (e.g., Purchase) Events . Finally, we incorporate the purchase events as prompts, such as a user has interacted with (purchased) ¡ITEM LIST¿.
In this way, we aim to augment the prompt texts and enrich the contextual environment that simulates the preference, characteristics, functionality, categorization, opinion, and first-order and second-order social or network dimensions of a real recommender system in a language modality. These crowd contextual prompts are used as training data to pre-train our graph-attentive LLM.
Figure 2. The structure of crowd contextual prompts used for pre-training. The prompts include: (1) User/Item Tokens to encode semantic information about users and items; (2) Item Contents, describing item attributes such as title, brand, and categories; (3) 1st Order User-Item Relationships, capturing user interactions through reviews and purchase explanations; (4) 2nd Order User-Item Relationships, representing item-level associations; and (5) User-Item Interaction Events, incorporating purchase histories as prompts.
@@ 3.3.3. The Optimization Objective and Procedures of Pre-training
(5) $\mathcal{L}_{\text{pre-train}}=-\sum_{i}\log P(t_{i+1}|t_{i},t_{i-1},\ldots,t_{1};\Theta),$
where $t_{i},t_{i-1},\ldots,t_{1}$ is the first $i$ tokens, and $\Theta$ indicates the weights of the LLM, the objective is to predict the $i+1$ token $t_{i+1}$. By solving the optimization objective, the graph-attentive LLM can combine knowledge from different information sources to learn more comprehensive portraits of users and items. By integrating graph-structured attention, the graph attentive LLM can model the first-order and second-order edge connectivity among users and items.
@@ 3.4. Fine-tuning Graph Attentive LLM with Personalized Predictive Prompts
After the pre-training step, the graph-attentive LLM learns the contextual knowledge of users, items, and relationships of a recommender system’s world. However, the optimization objective in the pre-training step focuses on maximizing textual generation accuracy in a language sequence instead of recommending personalized items. Therefore, in the fine-tuning step, we develop (1) personalized predictive prompts and (2) recommendation loss functions of fine-tuning to incentivize the graph-attentive LLM to shift model focuses from text generation to generating accurate personalized item recommendations.
@@ 3.4.1. Personalized Predictive Prompts
When fine-tuning the pre-trained graph attentive LLM, we introduce the personalized predictive Prompt method. Our idea is to use the historical purchase events of users for items as prompts to guide the LLM to learn user preferences for items. Fig. 3 shows that we convert a user’s interaction (e.g., rating, purchases) history with all items into past tense texts, combined with a future tense trigger (e.g., user $i$ will purchase ????), to motivate the graph attentive LLM to generate item recommendations for a user.
Figure 3. The illustration of the Personalized Predictive Prompt used for fine-tuning. The prompt converts a user’s past interactions into structured past-tense texts, followed by a future-tense trigger. This design encourages the LLM to infer user preferences and generate personalized item recommendations.
@@ 3.4.2. The Optimization Objective and Training Procedure of Fine-tuning
Different from learning the world knowledge of a recommender system, the fine-tuning stage is to adapt the pre-trained LLM to personalized item recommendations. During the fine-tuning process, we will integrate personalized predictive prompts into LLM so that the model can model the historical item purchase event of a user to generate a list of items $\mathcal{S}_{i}$ to the specific user, based on the prompt. The generative recommended items are then compared with the actual purchase records of the specific user. The resulting loss function derived from this comparison is utilized as the optimization objective of fine-tuning. In particular, we define the generative probability that measures whether LLM recommendations are statistically close to historical user purchase records, which is given by Equation (6):
(6) $\mathcal{L}_{\text{fine-tune}}=-\sum_{i}\log P(\mathcal{S}_{i}|\text{Pr}_{i},\Theta),$
where $\text{Pr}_{i}$ is the prompt for fine-tuning stage, and $\mathcal{S}_{i}$ is the final recommendation list, $\Theta$ denotes the LLM weights. In summary, fine-tuning optimizes the recommendation loss.
@@ 3.5. Graph Attentive LLM for Item Recommendations in Deployment
After pre-training and fine-tuning, given a testing user $i$, we convert the user’s interaction records with items into a personalized predictive prompt as the input of the graph-attentive LLM. Therefore, the classic recommendation engine in production can be viewed as a wrapper, where the request is reconstructed as a prompt -the same as the prompt in the fine-tuning stage- and the graph-attentive LLM will provide a recommendation score for each user-item pair.
To reduce the serving latency in production and alleviate the peak load pressure, the proposed graph attentive LLM model can be deployed onto both offline and online GPU clusters. The offline pipeline focuses on the batch processing and calculates the relevance between a user and the candidate items to recommend, shown as $\mathbf{LLM}(\text{Pr}_{i};\Theta)$, where $\text{Pr}_{i}$ is the user-item interaction prompt of user $i$, and $\Theta$ denotes the weights for the $\mathbf{LLM}$ backbone. The batch processing can be applied directly in offline recommendation scenarios such as promotional emails and notifications. The results can also be used for warming up the online cache to minimize redundant computations.
In a typical online recommendation scenario, the prompt containing the user, item, and interaction information is sent to the graph attentive LLM model, and the top items with the highest scores are selected as recommendations by comparing the probability scores against each other,
@@
In this case, the proposed method can easily be integrated into the most common recommender systems in the industry. The additional pressure caused by GPU serving can be handled by offline (batch) pre-computation and online caching warm-up.
@@ 4. Experimental Results
We conduct empirical experiments to answer the following questions: (1) Can our method generate more accurate recommendations? (2) What are the contributions of different technical components? (3) What are the contributions of second-order relationships and item background information? (4) What are the impacts of different attention mechanisms? (5) parameter sensitivity and robustness.
@@ 4.1.1. Data Description
We used seven public recommendation datasets: Amazon (AM)-Beauty dataset, AM-Toys dataset, AM-Sports, AM-Luxury, AM-Scientific, and AM-Instruments dataset (McAuley and Yang, 2016). We binarized the user-item interaction matrix by scores. If the score is greater than 3, there is a connection between a user and an item. For each user in the dataset, we randomly select 80% of interactions for training, 10% for validation, and 10% for testing, with at least one sample selected in both the validation and test sets. According to the prompt construction method in Section 3.3, we constructed the data for pre-training. Table 2 shows The main dataset statistics.
@@ 4.1.3. Baseline Algorithms
We compared our method with various approaches, including ID-based and Attention-based methods:
  • Multi-VAE (Liang et al., 2018) is an ID-based collaborative filtering method that completes recommendation tasks by using a polynomial likelihood variational auto-encoder to reconstruct ratings.
@@
  • $S^{3}Rec$ (Zhou et al., 2020) extends BERT4Rec by adding auxiliary tasks, such as item attribute prediction to enhance MLM, which can integrate content features for self-supervised learning.
@@
  • LightGCN (He et al., 2020) ignores feature transformation and nonlinear activation to enhance training efficiency and recommendation performance.
  • DWSRec (Zhang et al., 2024) is a sequential recommendation method that relies solely on pre-trained text embeddings and introduces a dual-view whitening strategy to enhance their effectiveness.
  • HSTU (Zhai et al., 2024) is a generative recommendation architecture that reformulates recommendation as a sequential transduction task, achieving superior accuracy and scalability on large-scale, high-cardinality data and demonstrating power-law scaling similar to foundation models like GPT-3.
  • LLMRec (Wei et al., 2024) is a graph augmentation framework that uses large language models to enrich user-item interaction graphs through edge reinforcement, item attribute enhancement, and user profiling.
  • RecMind (Wang et al., 2023) is an LLM-powered autonomous recommender agent that performs zero-shot personalized recommendation by planning with external tools and knowledge.
@@ 4.1.4. Hyperparameters and Settings
We conducted experiments using GPT-2 as the base model. We set the maximum input length to 1024, the token embedding dimension to 768, and the vocabulary length of natural language tokens to 50257. For Equation (4), $\delta$ is set to 0.9. In the pre-training stage, we first trained 10 epochs using crowd contextual data to optimize LLM and then trained 100 rounds using user-item interaction data. In the fine-tuning stage, we used 50 epochs for the recommendation-oriented fine-tuning of LLM.
@@ 4.1.5. Experimental Environment
All experiments were conducted on Ubuntu 22.04.3 LTS OS, Intel(R) Core(TM) i9-13900KF CPU, with the framework of Python 3.11.5 and PyTorch 2.0.1. All data are computed on an NVIDIA GeForce RTX 4090 GPU, which features 24,576 MiB of memory with CUDA version 12.2.
@@ 4.2.1. Overall Comparison
This experiment aims to answer: Can our model really generate more accurate recommendation results through the natural language processing method? We compared our model with several baseline models on various Amazon datasets. The baseline models used for comparison include ID-based and Attention-based methods. Our model was tested on the same dataset as these baseline models to ensure fairness and accuracy in the comparison. The experimental results are shown in Table 3, and our model performs well in seven Amazon datasets. Recall@20, Recall@40, and NDCG@100 are superior to the baseline models. This indicates that LLMs have strong capabilities in understanding text and capturing user preferences and needs, thereby promoting the accuracy of recommendations. Overall, the experimental results support our hypothesis that our model can generate more accurate recommendations through the graph-attentive LLM. This discovery is important for research and the practical application of recommender systems.
We also observe that some baselines perform better than others on certain datasets. For example, methods like UniSRec and FDSA perform well on datasets such as AM-Luxury and AM-Scientific. These models benefit from rich sequential user interactions and are optimized for capturing fine-grained temporal or semantic patterns. In contrast, traditional collaborative filtering methods (e.g., Multi-VAE, LightGCN) tend to underperform on such datasets due to their limited capacity for encoding semantic context. Our model shows consistent improvements across datasets by leveraging both textual semantics and graph connectivity in a unified prompt-driven manner.
In addition, we found that some models, including ours, are more sensitive to the characteristics of different datasets. For instance, AM-Food and AM-Instruments are relatively sparse and contain shorter user sequences. Models like SASRec and GRU4Rec, which depend on longer sequences, are affected more on these datasets. Our model maintains strong performance in these cases, as it integrates higher-order user-item relationships via graph structure and uses prompt-based representations to compensate for interaction sparsity. These observations highlight the importance of incorporating both structure and semantics in low-data or sparse interaction scenarios.
@@
Dataset Metric Multi-VAE MD-CVAE LightGCN BERT4Rec $S^{3}$ec UniSRec FDSA SASRec GRU4Rec DWSRec HSTU LLMRec RecMind Ours
AM-Beauty Recall@20 0.1295 0.1472 0.1429 0.1126 0.1354 0.1462 0.1447 0.1503 0.0997 0.1510 0.1546 0.1508 0.1347 0.1590
Recall@40 0.1720 0.2058 0.1967 0.1677 0.1789 0.1898 0.1875 0.2018 0.1528 0.1985 0.2104 0.2018 0.1874 0.2177
NDCG@100 0.0835 0.0871 0.0890 0.0781 0.0867 0.0907 0.0834 0.0929 0.0749 0.0971 0.0973 0.0927 0.0846 0.1029
AM-Toys Recall@20 0.1076 0.1107 0.1096 0.0853 0.1064 0.1110 0.0972 0.0869 0.0657 0.1307 0.912 0.1207 0.1126 0.1349
Recall@40 0.1558 0.1678 0.1558 0.1375 0.1524 0.1457 0.1268 0.1146 0.0917 0.1749 0.1208 0.1639 0.1564 0.1873
NDCG@100 0.0781 0.0812 0.0775 0.0532 0.0665 0.0638 0.0662 0.0525 0.0439 0.0784 0.0569 0.0672 0.0584 0.0876
AM-Sports Recall@20 0.0659 0.0714 0.0677 0.0521 0.0616 0.0714 0.0681 0.0541 0.0720 0.0753 0.0631 0.0701 0.0683 0.0764
Recall@40 0.0975 0.1180 0.0973 0.0701 0.0813 0.1143 0.0866 0.0739 0.1086 0.0960 0.0868 0.1183 0.1147 0.1240
NDCG@100 0.0446 0.0514 0.0475 0.0305 0.0438 0.0504 0.0475 0.0361 0.0498 0.0484 0.0399 0.0498 0.0511 0.0535
AM-Luxury Recall@20 0.2306 0.2771 0.2514 0.2076 0.2241 0.3091 0.2759 0.2550 0.2126 0.2524 0.2779 0.2761 0.2879 0.3066
Recall@40 0.2724 0.3206 0.3004 0.2404 0.2672 0.3675 0.3176 0.3008 0.2522 0.2876 0.3206 0.3219 0.3351 0.3441
NDCG@100 0.1697 0.2064 0.1947 0.1617 0.1542 0.2010 0.2107 0.1965 0.1623 0.1476 0.2177 0.2017 0.2049 0.2331
AM-Scientific Recall@20 0.1069 0.1389 0.1385 0.0871 0.1089 0.1492 0.1188 0.1298 0.0849 0.1096 0.1401 0.1409 0.1274 0.1480
Recall@40 0.1483 0.1842 0.1857 0.1160 0.1541 0.1954 0.1547 0.1776 0.1204 0.1360 0.1748 0.1839 0.1651 0.1908
NDCG@100 0.0766 0.0872 0.0834 0.0606 0.0715 0.1056 0.0846 0.0864 0.0594 0.0645 0.0927 0.0978 0.0873 0.1072
AM-Instruments Recall@20 0.1096 0.1398 0.1195 0.1183 0.1352 0.1684 0.1382 0.1483 0.1271 0.1057 0.1563 0.1322 0.1539 0.1698
Recall@40 0.1628 0.1743 0.1575 0.1531 0.1767 0.2239 0.1787 0.1935 0.1660 0.1423 0.2103 0.1727 0.2093 0.2265
NDCG@100 0.0735 0.1040 0.0985 0.0922 0.0894 0.1075 0.1080 0.0934 0.0998 0.0682 0.1074 0.0937 0.1008 0.1312
AM-Food Recall@20 0.1062 0.1170 0.1149 0.1036 0.1157 0.1423 0.1099 0.1171 0.1140 0.1341 0.1204 0.1347 0.1194 0.1438
Recall@40 0.1317 0.1431 0.1385 0.1284 0.1456 0.1661 0.1317 0.1404 0.1389 0.1618 0.1477 0.1564 0.1399 0.1673
NDCG@100 0.0727 0.0863 0.0853 0.0835 0.0926 0.1024 0.0904 0.0942 0.0910 0.0823 0.0929 0.0993 0.0783 0.1119